OpenStudy (anonymous):

Factor completely. If a polynomial is prime, state this. 9t^2 + 14t + 5

OpenStudy (anonymous):

(1 + t)(5 + 9t)

OpenStudy (anonymous):

lets see if i understand "factor by grouping" take \[9\times 5=45\] then i want two numbers whose product is 45 and whose sum is 14

OpenStudy (anonymous):

those would be 9 and 5 because \[9+5=14\] and \[9\times 5=45\]

OpenStudy (dboyette):

9t^(2)+14t+5 Find the factors such that the product of the factors is the trinomial 9t^(2)+14t+5. This can be done by trial and error and checked using the FOIL method of simplifying polynomials. (t+1)(9t+5)

OpenStudy (anonymous):

then write \[9t^2+9t+5t+5\] \[9t(t+1)+5(t+1)=(9t+5)(t+1)\]

OpenStudy (anonymous):

First get the product of 9 x 5 which gives us 45 Now find two positive factors of 45 whose sum is 14. 9 x 5 is our factors and the sum of 9 + 5 is 14 so now we use the grouping method 9t^2 + 14t + 5 9t^2 + 9t+ 5t + 5 Now Group (9t^2 + 9t)/9t+ (5t + 5)/5 9t(t+1)+ 5(t+1) = (9t+5)(t+1) Man I am slow at responding lol

OpenStudy (anonymous):

same answer as what I had